# Topological phase transitions in perovskite superlattices driven by temperature, electric field, and doping

### v2

@@ III.5 Temperature-driven phase transitions in (PTO)10/(PSTO)10 superlattices
We select a representative composition of PSTO with $\alpha=0.4$ to explore the temperature-dependent phase transitions in (PTO)10/(PSTO)10 superlattices, using 40$\times$20$\times$20 supercells at a strain state of $\delta_{\rm ave}=0.6$% (Fig. 6a). The Pb doping recovers the out-of-plane polarization ($P_{z}$), which generally decreases with increasing temperature. The system undergoes a conventional ferroelectric-paraelectric phase transition along the out-of-plane direction at $T\approx 550$ K (see $P_{z}$ versus $T$ in Fig. 6b). Similar to the FE-AFE-PE successive phase transitions observed in the (PTO)10/(STO)10 superlattice along the $x$ axis, the (PTO)10/(PSTO)10 superlattice maintains this transition sequence (see $P_{x}$ versus $T$ in Fig. 6b). As shown in Fig. 6c, the zigzag ordering of vortex cores in the $xz$ plane at 100 K results in a net $P_{x}$, which evolves to a horizontally aligned state at 300 K (Fig. 6d) associated with zero $P_{x}$ components. We note that although the spacial distributions of dipoles in the $xz$ plane resemble polar vortex arrays at both 100 K and 300 K, detailed three-dimensional dipole distribution analysis confirms that the dipoles form enlarged skyrmion bubbles embedded in $c$ domains (see Fig. 6e-f).
By comparing with standard skyrmion bubble formed in undoped (PTO)10/(STO)10 grown on an STO substrate ($\delta_{\rm ave}=0.0$), we identify several distinct features of the enlarged skyrmion bubbles. The dipole distributions in the $xy$ plane of standard skyrmion bubbles (Fig. 6g) reveal sharp boundaries and a low volume fraction of pure $c$ domains. In contrast, at 100 K, the skyrmion bubble residing in $c$ domains features diffusive boundaries (Fig. 6f). The bubble is anisotropic, consistent with the non-zero $P_{y}$ component shown in Fig. 6b. At 300 K, the enlarged bubble becomes more isotropic (Fig. 6f), associated with a zero net $P_{y}$.
@@ III.7 Concerted effects of doping, strain, and lattice periodicity
Doping strategies can also be combined with epitaxial strain engineering and superlattice periodicity control to realize diverse polar topological structures. Previously, it has been established experimentally that epitaxial strain engineering can be employed to achieve skyrmion bubbles in (PTO)16/(STO)16 grown on a STO substrate (with $\delta_{\mathrm{ave}}=0.0$%) as well as polar vortex arrays in (PTO)10/(STO)10 grown on a DSO substrate (with $\delta_{\mathrm{ave}}=1.0$%). This trend is reproduced by our MD simulations here (Fig.  8a-b). As discussed above, the introduction of Pb dopants into the STO layers at $\delta_{\mathrm{ave}}\approx 0.60$% can drive the superlattices from the vortex state to a mixed state with enlarged skyrmion bubbles and $c$ domains (Fig. 8b). Interestingly, further applying an in-plane tensile strain to $\delta_{\mathrm{ave}}\approx 0.70$% will recover the vortex arrays featuring a larger distance between vortex cores, as shown in Fig. 8d. An even more complex topology can be achieved by simultaneously adjusting the strain, doping concentration, and the lattice periodicity. For example, as recently demonstrated in our work, the (PTO)16/(Pb0.5Sr0.5TiO3)20 superlattice at $\delta_{\mathrm{ave}}\approx 1.0$% supports vortex arrays in PTO layers while forming dipole spirals in the Pb0.5Sr0.5TiO3 layers [42]. Large-scale MD simulations at finite temperatures, enabled by accurate deep potential capable of describing superlattices of different compositions, thus offer exciting opportunities to discover new types of ferroelectric topological structures and their dynamical properties.
@@ IV Conclusions
Our large-scale molecular dynamical simulations, enabled by accurate machine-learning force fields, demonstrate that the topologically engineered (PbTiO3)10/(SrTiO3)10 superlattice exhibits a range of unique characteristics along the in-plane direction. These include a low coercive field, distinctive phase transitions, and promising potential for tunable dielectric applications. A particularly intriguing finding is the counterintuitive temperature-driven transition from a ferroelectric-like to an antiferroelectric-like state. This unexpected behavior underscores the potential of harnessing topological domain structures to engineer properties that are unattainable in conventional single-domain states. The regulation of the depolarization field in the PbTiO3 layers through strategic doping of the SrTiO3 layers is effective in driving dipolar topological phase transitions and fostering the formation of new types of complex polar topologies. The quantitative atomistic insights into the thermal and electrical responses of these perovskite superlattices provide a foundation for tailoring their properties for nanoelectronic applications.

### v4

@@ III.5 Temperature-driven phase transitions in (PTO)10/(PSTO)10 superlattices
We select a representative composition of PSTO with $\alpha=0.4$ to explore the temperature-dependent phase transitions in (PTO)10/(PSTO)10 superlattices, using 40$\times$20$\times$20 supercells at a strain state of $\delta_{x}=0.46$% and $\delta_{y}=0.26$% (Fig. 8a). The Pb doping recovers the out-of-plane polarization ($P_{z}$), which generally decreases with increasing temperature. The system undergoes a conventional ferroelectric-paraelectric phase transition along the out-of-plane direction at $T\approx 550$ K (see $P_{z}$ versus $T$ in Fig. 8b). Similar to the FE-AFE-PE successive phase transitions observed in the (PTO)10/(STO)10 superlattice along the $x$ axis, the (PTO)10/(PSTO)10 superlattice maintains this transition sequence (see $P_{x}$ versus $T$ in Fig. 8b). As shown in Fig. 8c, the zigzag ordering of vortex cores in the $xz$ plane at 100 K results in a net $P_{x}$, which evolves to a horizontally aligned state at 300 K (Fig. 8d) associated with zero $P_{x}$ components. We note that although the spacial distributions of dipoles in the $xz$ plane resemble polar vortex arrays at both 100 K and 300 K, detailed three-dimensional dipole distribution analysis confirms that the dipoles form enlarged skyrmion bubbles embedded in $c$ domains (see Fig. 8e-f).
By comparing with standard skyrmion bubble formed in undoped (PTO)10/(STO)10 grown on an STO substrate ($\delta_{x}=$0.23 %, $\delta_{y}=$0.23 %), we identify several distinct features of the enlarged skyrmion bubbles. The dipole distributions in the $xy$ plane of standard skyrmion bubbles (Fig. 8g) reveal sharp boundaries and a low volume fraction of pure $c$ domains. In contrast, at 100 K, the skyrmion bubble residing in $c$ domains features diffusive boundaries (Fig. 8e). The bubble is anisotropic, consistent with the non-zero $P_{y}$ component shown in Fig. 8b. At 300 K, the enlarged bubble becomes more isotropic (Fig. 8f), associated with a zero net $P_{y}$.
@@ III.7 Concerted effects of doping, strain, and lattice periodicity
Doping strategies can also be combined with epitaxial strain engineering and superlattice periodicity control to realize diverse polar topological structures. Previously, it has been established experimentally that epitaxial strain engineering can be employed to achieve skyrmion bubbles in (PTO)16/(STO)16 grown on a STO substrate with a slightly compressive strain [7], as well as polar vortex arrays in (PTO)10/(STO)10 grown on a DyScO3 substrate with a tensile strain [6]. This trend is reproduced by our MD simulations here (Fig.  10a-c). As discussed above, the introduction of Pb dopants into the STO layers at $\delta_{x}=0.46$% and $\delta_{y}=0.26$% can drive the superlattices from the vortex state to a mixed state with enlarged skyrmion bubbles and $c$ domains (Fig. 10b). Interestingly, further applying an in-plane tensile strain to $\delta_{x}=0.61$% and $\delta_{y}=0.26$% will recover the vortex arrays featuring a larger distance between vortex cores, as shown in Fig. 10d. An even more complex topology can be achieved by simultaneously adjusting the strain, doping concentration, and the lattice periodicity. For example, as recently demonstrated in our work, the (PTO)16/(Pb0.5Sr0.5TiO3)20 superlattice at $\delta_{x}=0.84$% and $\delta_{y}=0.69$% supports vortex arrays in PTO layers while forming dipole spirals in the Pb0.5Sr0.5TiO3 layers [56]. Large-scale MD simulations at finite temperatures, enabled by accurate deep potential capable of describing superlattices of different compositions, thus offer exciting opportunities to discover new types of ferroelectric topological structures and their dynamical properties.
@@ IV Conclusions
Our large-scale molecular dynamical simulations, enabled by accurate machine-learning force fields, demonstrate that the topologically engineered (PbTiO3)10/(SrTiO3)10 superlattice exhibits a range of unique characteristics along the in-plane direction. These include a low coercive field, distinctive phase transitions, and promising potential for tunable dielectric applications. A particularly intriguing finding is the counterintuitive temperature-driven transition from a ferroelectric-like (ferrielectric-like) to an antiferroelectric-like state. This unexpected behavior underscores the potential of harnessing topological domain structures to engineer properties that are unattainable in conventional single-domain states. The regulation of the depolarization field in the PbTiO3 layers through strategic doping of the SrTiO3 layers is effective in driving dipolar topological phase transitions and fostering the formation of new types of complex polar topologies. The quantitative atomistic insights into the thermal and electrical responses of these perovskite superlattices provide a foundation for tailoring their properties for nanoelectronic applications.